The oxidation number of arsenic (As) in NaAsO, is +5
Answer #1

Oxidation number of Na = +1

Oxidation number of O = -2

lets the oxidation number of As be x

use:

1* oxidation number (Na) + 2* oxidation number (O) + 1* oxidation number (As) = net charge

1*(+1)+2*(-2)+1* x = 0

-3 + 1 * x = 0

x = 3

So oxidation number of As = +3

Answer: +3